Show that the test taking the overall significance of regression model using ANOVA table to be expressed as:  Where, R be a level of determination and k is the number of parameters in the n sampled regression model


1
Expert's answer
2021-10-11T10:44:28-0400

"F=\frac {MSR} {MSE}"


Here, MSR is the mean squared anomaly and MSE is the mean squared error.


"MSR=\frac {SSR} {1}" , called regression mean square


"MSE=\frac {SSE} {(n-2)}" called error mean square.


"F=\frac {1-R^2} {n-k}"